In eight hours A walks twelve km more then B does in seven hours; and in thirteen hours B walks seven km more then A does in nine hours. How many km does each walk per hour?
1
Expert's answer
2012-09-07T10:53:46-0400
8A-7B=12 and 13B-9A=7 adding this equations we get 6B-A=19 so A=6B-19 than
48B-152-7B=12 41B=164 B=4 A=5